OpenStudy (anonymous):

Assuming 100% dissociation, calculate the freezing point and boiling point of 2.59 m K3PO4(aq).

OpenStudy (anonymous):

Use the formula T=Kf*i*m Here i=4 as K3PO4 disociates into 4 ions. Kf=1.86 so you get T=4*1.86*2.59=19.2696 Hence freezing point=0-19.2696=-19.2696 degree Celsius